# Unveiling Client Privacy Leakage from Public Dataset Usage in Federated Distillation

### v1

@@ Abstract.
Federated Distillation (FD) has emerged as a popular federated training framework, enabling clients to collaboratively train models without sharing private data. Public Dataset-Assisted Federated Distillation (PDA-FD), which leverages public datasets for knowledge sharing, has become widely adopted. Although PDA-FD enhances privacy compared to traditional Federated Learning, we demonstrate that the use of public datasets still poses significant privacy risks to clients’ private training data. This paper presents the first comprehensive privacy analysis of PDA-FD in presence of an honest-but-curious server. We show that the server can exploit clients’ inference results on public datasets to extract two critical types of private information: label distributions and membership information of the private training dataset. To quantify these vulnerabilities, we introduce two novel attacks specifically designed for the PDA-FD setting: a label distribution inference attack and innovative membership inference methods based on Likelihood Ratio Attack (LiRA). Through extensive evaluation of three representative PDA-FD frameworks (FedMD, DS-FL, and Cronus), our attacks achieve state-of-the-art performance, with label distribution attacks reaching minimal KL-divergence and membership inference attacks maintaining high True Positive Rates under low False Positive Rate constraints. Our findings reveal significant privacy risks in current PDA-FD frameworks and emphasize the need for more robust privacy protection mechanisms in collaborative learning systems.
@@ 1. Introduction
In this work, we devise two types of privacy attacks against clients by the server: Label Distribution Inference Attacks (LDIA) and Membership Inference Attacks (MIA). LDIA reveals privacy information about the overall data distribution of all client private datasets combined. In contrast, membership inference attacks enable more granular privacy leakage by identifying the presence of specific data samples in a client’s private datasets. Both LDIA and MIA can be performed in a black-box manner, requiring only clients’ inference results on the public dataset, which renders these attacks practically useful since FD frameworks are generally designed to not transfer a client model to the server. Additionally, Federated Distillation assumes non-IID data distributions across clients in the label space. LDIA becomes particularly important in this context, as inferring the label distribution can reveal significant information about a client’s unique data characteristics. Furthermore, LDIA and MIA can serve as stepping stones for more sophisticated attacks. For instance, the obtained knowledge can be leveraged to generate synthetic data that mimics private datasets.
@@
In the recent literature, LDIA and MIA have been extensively studied in the traditional FL and centralized machine learning settings (Nasr et al., 2019; Shokri et al., 2017; Yeom et al., 2018; Melis et al., 2019; Ramakrishna and Dán, 2022; Jiang et al., 2024). They have largely focused on white-box or gradient-based attacks. For example, Gu et al. demonstrated that a malicious server in FL can infer label distributions by exploiting the gradients or parameters uploaded by clients (Gu and Bai, 2023). Similarly, Wainakh et al. showed that user-level label leakage is possible through gradient analysis (Wainakh et al., 2021). For MIAs, Nasr et al. developed sophisticated MIA techniques that exploit the white-box access to model parameters in FL (Nasr et al., 2019). The rich information leveraged in these attacks are not directly accessible in FD settings. In the FD setting, a few studies have primarily focused on MIAs, such as FD-Leaks (Yang et al., 2022), MIA-FedDL (Liu and Dong, 2023) and GradDiff (Wang et al., 2024). In Fed-leaks, a malicious client leverages its local model as a shadow model to train an MIA classifier, then applies the classifier on target sample inference score to determine target sample membership in other clients’ private datasets. The attacker needs to be the client in the  MIA-FedDL approach;  GradDiff can be used by either a client or the server to attack another client. Both of these approaches require shadow datasets to train shadow models for MIA. However, effective MIA requires the attacker’s shadow model to be trained on the dataset sharing similar data distribution with the target model’s training dataset. In the FD setting, it is challenging for both clients and the server to obtain the data distribution of other clients’ private data. Therefore, these MIA methods cannot work effectively in heterogeneous non-IID environments.
@@
To address the limitations of the existing works, we aim to comprehensively examine privacy leakage by public datasets in FD across multiple frameworks (FedMD, DS-FL, and Cronus) and various data distribution scenarios. We also introduce new attack methods that are specifically tailored to the PDA-FD setting. Specifically, we propose a novel LDIA method based on public datasets and extend the state-of-the-art MIA, Likelihood Ratio Attack (LiRA) (Carlini et al., 2022), to overcome the challenges posed by the limited information available in PDA-FD. To that end, we design and implement Co-op LiRA and Distillation-based LiRA for MIAs. Furthermore, while previous works primarily focused on client-side attacks, our work provides a more holistic view by examining both LDIA and MIA from the server’s perspective. A high-level comparison with the existing work is shown in Table 1.
@@
Method Attacker Framework Shadow dataset Attack Goal
@@
2. Background
2.1. Federated Distillation
@@
In our study, we focus on one category of FD frameworks, Public Dataset-Assisted Federated Distillation (Li and Wang, 2019; Itahara et al., 2021; Chang et al., 2019)(PDA-FD). The workflow of PDA-FD is shown in Figure 1. In PDA-FD, the server leverages a public dataset to facilitate knowledge transfer among clients. The public dataset is shared with all the clients. The PDA-FD framework typically involves three phases in each collaborative training round: local updates phase, communication phase, and knowledge distillation phase.
During the local updates phase, client $n$ trains its local model $\theta_{n}$ on its private dataset $D_{n}$ using stochastic gradient descent (LeCun et al., 1998). The loss function $\mathcal{L}(x,y,\theta_{n})$ is defined to calculate the error between the prediction posterior $f_{\theta_{n}}(x)_{y}$ of the training data and its ground truth label $y$. Cross-entropy is often used as the loss function:
(1) $\mathcal{L}(x,y,\theta_{n})=-log(f_{\theta_{n}}(x)_{y})$
@@
During the knowledge distillation phase, client $n$ performs knowledge distillation (Hinton et al., 2015) using the aggregated logits $Z_{k}$ returned by the server as the soft labels. In this case, the loss function also needs to include the mean absolute errors (MAE):
@@
2.2. Label Distribution Inference Attack
In both FL and FD, Label Distribution Inference Attacks (LDIA) can pose significant threats to individual clients’ privacy. In an LDIA, the adversary aims to infer the proportion of training data across different labels in the target client’s training dataset. Previous work (Gu and Bai, 2023; Wainakh et al., 2021) has demonstrated that the server of FL can infer the distribution of clients’ training data by exploiting the gradients or parameters uploaded by clients in FL, leading to privacy leakage.
Definitions. Given a target client’s local model $\theta_{n}$ that is trained on a private dataset $D_{n}$, which consists of data samples from $M$ classes: $D_{n}=\bigcup_{m=1}^{M}D_{n}^{m}$, where $D_{n}^{m}$ refers to the subset containing data of label m. The ground truth label distribution of $D_{n}$ is calculated as follows:
(4) $\mathbf{p}=\left(p_{1},p_{2},\dots,p_{M}\right),\quad p_{m}=\frac{|D_{n}^{m}|}{|D_{n}|}$
where $p_{m}$ represents the proportion of data with label $m$ in $D_{n}$, and $\sum_{m=1}^{M}p_{m}=1$ and $p_{m}\in[0,1]$. The objective of LDIA can be expressed as a function $\mathcal{A}$ that maps from the observable information about the target model outputs to an estimated label distribution:
(5) $\mathcal{A}:\theta_{n}\mapsto\mathbf{\hat{p}}=\left(\hat{p}_{1},\hat{p}_{2},\dots,\hat{p}_{M}\right)$
where $\hat{p}_{m}$ represents the inferred proportion of data with label $m$ in $D_{n}$.
The success of LDIA can compromise the privacy of FL and FD systems, as it allows attackers to gain insights into the composition of private datasets without direct access.
2.3. Membership Inference Attack
In membership inference attacks (MIA), the attacker’s objective is to determine whether a target data sample is in the target model’s training dataset. As machine learning models become more widely deployed, their training datasets may contain sensitive information (Havaei et al., 2017; Fu et al., 2016), and MIA poses a significant threat to the privacy of such data. FD, which was designed to protect data privacy, has also become the target of various MIA attacks (Yang et al., 2022; Wang et al., 2024) against clients’ local models within the framework.
Definitions. Given a target model $\theta$, and a target data sample $x$, the objective of MIA can be define as a function $\mathcal{A}$:
(6) $\mathcal{A}:x,\theta\mapsto\{0,1\}$
where the output is 1 if $x$ is inferred to be in the training dataset of model $\theta$, and 0 otherwise. In MIA, we refer to the data sample belonging to the target model’s training set as members, while other data samples are referred to as non-members. MIA exploits the behavioral differences of machine learning models on data they were trained on versus data they weren’t, leveraging the fact that models often exhibit higher confidence or lower loss on their training data due to overfitting.
@@ 3.1. Threat Model
Adversary Knowledge: We investigate PDA-FD in the context of Horizontal FL (HFL), where clients’ private data’s label distributions can be non-IID. We consider a threat model where the server acts as an honest-but-curious (Paverd et al., 2014) adversary. The server attacker is not allowed to modify the learning process but gets to select public dataset members used for knowledge transfer; this privilege for the server is common in PDA-FD frameworks (Li and Wang, 2019; Itahara et al., 2021; Chang et al., 2019). The server can select members of the public dataset in each collaborative training round. In each training round, the server’s access to clients’ models is black, and it only interacts with client models by running inferences on public data. The server can only obtain logit vectors or prediction vectors of every public data sample each client model provides without visibility into the private models’ weights or architectures.
Adversary’s objective:. The server aims to infer sensitive information about a target client’s private dataset, specifically label distribution information and membership information, through two main attack vectors: LDIA and MIA.
@@ 3.2. Label Distribution Inference Attack
Unlike traditional FL where the server has white-box access to clients’ local models, the server only has black-box access by using a public dataset in PDA-FD. The core idea behind LDIA is that the logits or prediction values produced by a client’s model still carry information about the distribution of labels in its training data. This is due to the tendency of neural networks to overfit their training distribution, even when regularization techniques are applied (Yeom et al., 2018).
@@
3.3. Membership Inference Attack
@@
While LiRA is a powerful and effective technique for MIAs, directly applying it in the context of PDA-FD presents challenges. In traditional LiRA, the attacker needs the ability to train multiple reference models that mimic the target model’s behavior but behave differently on the target samples, specifically, prediction discrepancies between target and reference models should be larger for members than for non-members. However, training such models requires access to a shadow dataset with a distribution similar to the target model’s training data (Watson et al., 2021; Carlini et al., 2022). As shown in Figure 5(a), when the data distribution of the reference model’s training dataset differs from that of the target model’s training dataset, the attacker cannot distinguish members based on prediction discrepancies between reference and target models for members versus non-members. The prediction discrepancy is quantified as the difference between the target model’s prediction probability and the mean of reference models’ prediction probabilities for the target sample. In the PDA-FD context, the server faces significant challenges in obtaining such a shadow dataset: (1) The FD framework allows clients to have non-IID private datasets in the label space. Without knowledge of the specific label distributions in each client’s private dataset, the server cannot accurately sample shadow datasets that match the characteristics of the target client’s data. (2) In FD, the public dataset can be unlabeled, particularly in semi-unsupervised learning scenarios. In that case, the server simply cannot use these datasets without labels for training shadow models, as label information is essential for mimicking the target model’s behaviors. To address these unique challenges, as shown in Figure 4, we design and implement two variants of offline LiRA: Co-op LiRA and Distillation-based LiRA.
@@
Figure 4. Workflow of Membership Inference Attack(Co-op LiRA and Distillation-based LiRA).

### v2

@@ Abstract.
Federated Distillation (FD) has emerged as a popular federated training framework, enabling clients to collaboratively train models without sharing private data. Public Dataset-Assisted Federated Distillation (PDA-FD), which leverages public datasets for knowledge sharing, has become widely adopted. Although PDA-FD enhances privacy compared to traditional Federated Learning, we demonstrate that the use of public datasets still poses significant privacy risks to clients’ private training data. This paper presents the first comprehensive privacy analysis of PDA-FD in the presence of an honest-but-curious server. We show that the server can exploit clients’ inference results on public datasets to extract two critical types of private information: label distributions and membership information of the private training dataset. To quantify these vulnerabilities, we introduce two novel attacks specifically designed for the PDA-FD setting: a label distribution inference attack and innovative membership inference methods based on Likelihood Ratio Attack (LiRA). Through extensive evaluation of three representative PDA-FD frameworks (FedMD, DS-FL, and Cronus), our attacks achieve state-of-the-art performance, with label distribution attacks reaching minimal KL-divergence and membership inference attacks maintaining high True Positive Rates under low False Positive Rate constraints. Our findings reveal significant privacy risks in current PDA-FD frameworks and emphasize the need for more robust privacy protection mechanisms in collaborative learning systems.
Federated Distillation, Membership Inference Attack, Label Distribution Inference Attack
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
@@ 1. Introduction
In this work, we devise two types of privacy attacks against clients by the server: Label Distribution Inference Attacks (LDIA) and Membership Inference Attacks (MIA). LDIA reveals privacy information about the overall data distribution of all client private datasets combined. In contrast, membership inference attacks enable more granular privacy leakage by identifying the presence of specific data samples in a client’s private datasets. Both LDIA and MIA can be performed in a black-box manner, requiring only clients’ inference results on the public dataset, which renders these attacks practically useful since FD frameworks are generally designed to not transfer a client model to the server. Additionally, Federated Distillation assumes non-IID data distributions across clients in the label space. LDIA becomes particularly important in this context, as inferring the label distribution can reveal significant information about a client’s unique data characteristics. For example, if hospitals were conducting FD to protect patient privacy, attackers could expose which facilities specialize in rare conditions, potentially de-anonymizing patients with unique medical profiles while also compromising institutional privacy. Furthermore, LDIA and MIA can serve as stepping stones for more sophisticated attacks. For instance, the obtained knowledge can be leveraged to generate synthetic data that mimics private datasets.
@@
In the recent literature, LDIA and MIA have been extensively studied in the traditional FL and centralized machine learning settings (Nasr et al., 2019; Shokri et al., 2017; Yeom et al., 2018; Melis et al., 2019; Ramakrishna and Dán, 2022; Jiang et al., 2024). They have largely focused on white-box or gradient-based attacks. For example, Gu et al. demonstrated that a malicious server in FL can infer label distributions by exploiting the gradients or parameters uploaded by clients (Gu and Bai, 2023). Similarly, Wainakh et al. showed that user-level label leakage is possible through gradient analysis (Wainakh et al., 2021). For MIAs, Nasr et al. developed sophisticated MIA techniques that exploit the white-box access to model parameters in FL (Nasr et al., 2019). The rich information leveraged in these attacks are not directly accessible in FD settings.
In the FD setting, a few studies have primarily focused on MIAs, such as FD-Leaks (Yang et al., 2022), MIA-FedDL (Liu and Dong, 2023) and GradDiff (Wang et al., 2024). However, due to reasons such as the use of low-quality shadow models or overly strong threat model assumptions, these methods tend to experience decreased effectiveness in real-world FD scenarios. In FD-leaks, a malicious client leverages its local model as a shadow model to train an MIA classifier, then applies the classifier on target sample inference score to determine target sample membership in other clients’ private datasets. Effective MIA requires attackers to possess high-quality shadow models, which are shadow models that can mimic target model performance. Such shadow models need to be trained on shadow datasets sharing similar data distributions with the target model’s training dataset (Shokri et al., 2017; Carlini et al., 2022). However, in FD-leaks, the distribution of the training dataset of the malicious client’s local model may not be consistent with the distribution of the target model’s training dataset, resulting in the shadow model used by the malicious client potentially being a low-quality shadow model, which leads to limited MIA performance. In MIA-FedDL (Liu and Dong, 2023), the attacker acts as clients, while in GradDiff (Wang et al., 2024), the attacker can be either clients or the server. However, for MIA-FedDL and GradDiff, they assume that the attacker can additionally obtain a carefully designed shadow dataset with a distribution consistent with the target client model’s training dataset distribution to train the shadow models used in MIA. We consider this to be an overly strong threat model assumption in the FD scenario. In the FD setting, it is challenging for both clients and the server to obtain the data distribution of other clients’ private data. Therefore, in a more realistic threat model where clients and servers cannot directly obtain the private data distribution of the target client, these MIA methods (Yang et al., 2022; Liu and Dong, 2023; Wang et al., 2024) in heterogeneous non-IID environments will experience decreased effectiveness due to the inability to use high-quality shadow models in MIA. Simultaneously, we note that these MIAs in FD scenarios still combine the use of low-precision classifier-based MIA (Shokri et al., 2017), resulting in their relatively limited performance on the key evaluation metric of True Positive Rate at low False Positive Rate Region (Carlini et al., 2022).
@@
To address the limitations of the existing works, we aim to comprehensively and effectively examine privacy leakage by public datasets in FD across multiple frameworks (FedMD, DS-FL, and Cronus) and various data distribution scenarios. We introduce new attack methods that are specifically tailored to the PDA-FD setting. Specifically, we propose a novel LDIA method based on public datasets and extend the state-of-the-art MIA, Likelihood Ratio Attack (LiRA) (Carlini et al., 2022), to overcome the challenges posed by the limited information available in PDA-FD. To that end, we design and implement Co-op LiRA and Distillation-based LiRA for MIAs. These two MIA approaches relax the constraint present in most current MIAs (Carlini et al., 2022; Watson et al., 2021; Shokri et al., 2017) that require the attacker to possess a shadow dataset with the same data distribution as the target model’s training dataset. In order to obtain shadow models that can effectively mimic the target model’s performance, in co-op LiRA, the server executes LDIA as a preliminary attack to obtain the label distribution of all clients’ private data, then selects client models with label distributions similar to the target client model’s as shadow models in MIA. In distillation-based LiRA, the server utilizes the public dataset used in the FD training process, using the target client model as a teacher model to distill several student models as shadow models in MIA.
Compared to previous MIA methods in FD, for FD-leaks (Yang et al., 2022), through our approach of selecting or training shadow models, we are able to obtain higher-quality shadow models that can better mimic the performance of target client models. Unlike MIA-FedDL (Liu and Dong, 2023) and GradDiff (Wang et al., 2024), our attack operates under more realistic threat model assumptions in FD scenarios, requiring only the existing public dataset from the FD training process rather than an additional carefully crafted shadow dataset that matches the target model’s training distribution. Furthermore, our proposed MIAs adapt and integrate the advanced MIA method LiRA into realistic FD scenarios with constrained threat model, offering significantly better performance than classifier-based MIAs (Shokri et al., 2017) used in prior works (Yang et al., 2022; Liu and Dong, 2023; Wang et al., 2024), thereby bringing privacy research in FD closer in line with the state-of-the-art developments in the MIA field (Carlini et al., 2022; Watson et al., 2021). Additionally, while some previous works (Yang et al., 2022; Wang et al., 2024) primarily focused on client-side attacks, our work provides a more holistic view by examining both LDIA and MIA from the server’s perspective. A high-level comparison with the existing work is shown in Table 1.
@@
Method Attacker Framework Ideal $D_{shadow}$* Attack Goal
@@
    $D_{shadow}$ that has the same data distribution as $D_{train}$.
@@
2. Understanding the Privacy Risks in PDA-FD
2.1. PDA-FD
@@
In our study, we focus on one category of FD frameworks, Public Dataset-Assisted Federated Distillation (PDA-FD) (Li and Wang, 2019; Itahara et al., 2021; Chang et al., 2019). The workflow of PDA-FD is shown in Figure 1. In PDA-FD, the server leverages a public dataset to facilitate knowledge transfer among clients. The public dataset is prepared by the server and shared with all the clients during the collaborative training rounds (Li and Wang, 2019; Chang et al., 2019; Itahara et al., 2021). The PDA-FD framework typically involves three phases in each collaborative training round: local updates phase, communication phase, and knowledge distillation phase.
During the local updates phase, client $n$ trains its local model $\theta_{n}$ on its private dataset $D_{n}$ using stochastic gradient descent (LeCun et al., 1998). The loss function $\mathcal{L}(x,y,\theta_{n})$ is defined to calculate the error between the prediction posterior $f_{\theta_{n}}(x)_{y}$ of the training data and its ground truth label $y$. Cross-entropy is often used as the loss function: $\mathcal{L}(x,y,\theta_{n})=-log(f_{\theta_{n}}(x)_{y})$.
@@
During the knowledge distillation phase, client $n$ performs knowledge distillation (Hinton et al., 2015) using the aggregated logits $Z_{k}$ returned by the server as the soft labels. These aggregated logits $Z_{k}$ represent knowledge learned by other clients’ private models on their own private datasets, allowing clients to learn from other clients’ knowledge and further improve their own model’s performance. In this case, the loss function will be the mean absolute errors (MAE):
@@
2.2. Privacy Risks Analysis
Federated Distillation (Chang et al., 2019; Li and Wang, 2019; Itahara et al., 2021) relies on clients and server using inference logits from public datasets to facilitate knowledge transfer and improve model performance. However, we find that in the existing PDA-FD algorithms (Chang et al., 2019; Li and Wang, 2019; Itahara et al., 2021), as shown in Algorithm 1, to better transfer knowledge from each client’s private model, it is often necessary to train the private model on the private dataset for few epochs during local updates phase before communication phase in collaborative training rounds. We believe that this behavior increases the private model’s overfitting level to its own private dataset, which inherently raises privacy leakage risks (Yeom et al., 2018).
Our analysis of FedMD training on CIFAR10 (Krizhevsky et al., 2009), as illustrated in Figure 2, confirms these concerns by tracking one client’s model accuracy and its overfitting to private dataset (measured by inference loss on private dataset) throughout collaborative training. Despite the initial training on both public and private datasets, model accuracy continues to improve significantly during collaborative rounds, demonstrating effective knowledge transfer via public datasets. However, during each round’s local updates(LU) phase, the overfitting level to private data substantially increases. More critically, during the subsequent communication phase, which occurs prior to the knowledge distillation(KD) phase, the server needs to interact with client private models through the public dataset to help process and distribute knowledge among all clients. This interaction, combined with the elevated overfitting level, makes privacy attacks feasible. Meanwhile, to ensure effective PDA-FD training and facilitate the processing of clients’ generated logits, PDA-FD protocols (Li and Wang, 2019; Chang et al., 2019; Itahara et al., 2021) typically require the server to prepare a public dataset for all the clients before training and to select subsets for knowledge distillation in each round. Therefore, an honest-but-curious server is capable of launching privacy attacks against a target client’s private dataset by carefully crafting the public dataset and leveraging the inference logits generated by the client. The privacy threat posed by such a server to clients should be given significant attention (Gu and Bai, 2023; Wang et al., 2024; Nasr et al., 2019).
Figure 2. Visualizing the client’s private model performance trend and overfitting trend to private dataset(members) through collaborative training rounds in FedMD (Li and Wang, 2019).
2.3. Training Data Privacy Attacks
Machine learning models face privacy attacks targeting training data. We focus on two key threats: Label Distribution Inference Attacks (LDIA) (Gu and Bai, 2023; Wainakh et al., 2021) and Membership Inference Attacks (MIA) (Shokri et al., 2017; Salem et al., 2020; Carlini et al., 2022; Liu et al., 2022; Ye et al., 2022), representing privacy violations on the dataset and the sample levels, respectively. They are also among the most investigated privacy attacks in the field. LDIA infers the proportion of different labels in a client’s private dataset. For a local model $\theta_{n}$ trained on dataset $D_{n}=\bigcup_{m=1}^{M}D_{n}^{m}$, where $D_{n}^{m}$ represents data of label $m$, the ground truth label distribution is: $\mathbf{p}=\left(p_{1},p_{2},\dots,p_{M}\right),\quad p_{m}=\frac{|D_{n}^{m}|}{|D_{n}|}$. The attack function maps from the observable information to an estimated label distribution: $\mathcal{A}:\theta_{n}\mapsto\mathbf{\hat{p}}=\left(\hat{p}_{1},\hat{p}_{2},\dots,\hat{p}_{M}\right)$.
In MIA, the attacker determines whether a data sample $(x,y)$ belongs to the target model $\theta$’s training dataset. This attack is defined as: $\mathcal{A}:x,\theta\mapsto\{0,1\}$, where 1 indicates member. MIA exploits the behavioral differences of models on training(member) versus non-training data(non-member), leveraging higher confidence or lower loss on training samples due to overfitting (Yeom et al., 2018; Shokri et al., 2017).
@@ 3.1. Threat Model
Adversary Knowledge. We investigate PDA-FD in the context of Horizontal FL (HFL), where clients’ private data’s label distributions can be non-IID. We consider a threat model where the server acts as an honest-but-curious (Paverd et al., 2014) adversary. The server attacker is not allowed to modify the learning process or affect the performance of FD training but gets to select public dataset members used for knowledge transfer; this privilege for the server is common in PDA-FD frameworks (Li and Wang, 2019; Itahara et al., 2021; Chang et al., 2019). The server can select members of the public dataset in each collaborative training round. In each training round, the server can only black-box access to clients’ models, and it only interacts with client models by running inferences on public data. The server can only obtain logit vectors or prediction vectors of every public data sample each client model provides without visibility into the private models’ weights or architectures.
Adversary’s Objective. The server aims to infer sensitive information about a target client’s private dataset, specifically label distribution information and membership information, through two main attack vectors: LDIA and MIA.
@@ 3.2. Label Distribution Inference Attack
In Federated Distillation, the server only has black-box access by using a public dataset in PDA-FD. The core idea behind our proposed LDIA is that the logits or prediction values produced by a client’s model still carry information about the distribution of labels in its training data. This is due to the tendency of neural networks to overfit their training distribution, even when regularization techniques are applied (Yeom et al., 2018).
@@
3.3. Membership Inference Attacks
@@
While LiRA is a powerful and effective technique for MIAs, directly applying it in the context of PDA-FD presents challenges. In traditional LiRA, the attacker needs the ability to train multiple reference models that mimic the target model’s behavior but behave differently on the target samples, specifically, prediction discrepancies between target and reference models should be larger for members than for non-members. However, training such models requires access to a shadow dataset with a distribution similar to the target model’s training data (Watson et al., 2021; Carlini et al., 2022). As shown in Figure 6(a), when the data distribution of the reference model’s training dataset differs from that of the target model’s training dataset, the attacker cannot distinguish members based on prediction discrepancies between reference and target models for members versus non-members. The prediction discrepancy is quantified as the difference between the target model’s prediction probability and the mean of reference models’ prediction probabilities for the target sample.
In the PDA-FD context, the server faces significant challenges in obtaining such a shadow dataset: (1) The FD framework allows clients to have non-IID private datasets in the label space. Without knowledge of the specific label distributions in each client’s private dataset, the server cannot accurately sample shadow datasets that match the characteristics of the target client’s data. (2) In FD, the public dataset can be unlabeled, particularly in semi-unsupervised learning scenarios. In that case, the server simply cannot use these datasets without labels for training shadow models, as label information is essential for mimicking the target model’s behaviors. To address these unique challenges, as shown in Figure 5, we design and implement two variants of offline LiRA: Co-op LiRA and Distillation-based LiRA. Section 5.5 also briefly explores how these approaches can be combined with indirect attack techniques (Long et al., 2020; Wen et al., 2022), further demonstrating the robustness of our proposed attacks.
@@
Figure 5. Workflows of two proposed Membership Inference Attacks(Co-op LiRA and Distillation-based LiRA).
3.3.1. Co-op LiRA
Our proposed Co-op LiRA introduces a novel approach that leverages our LDIA method as a prerequisite to enable effective membership inference attacks. Unlike traditional MIA techniques that require shadow datasets with similar distributions to the target, our method uniquely exploits the collaborative nature of federated distillation environments. The key innovation of Co-op LiRA lies in its ability to identify suitable reference models without additional training. We aim for these reference models to have training datasets with data distributions closer to the target model’s distribution, thereby better mimicking the target client model’s behavior (Shokri et al., 2017). First, the server conducts LDIA (as described in Section 3.2) to determine label distributions across all clients. In the HFL setting, clients with similar label distributions likely possess closer overall data distributions. This is because the data distribution may differ in two aspects: the feature and label space. In the definition of HFL, clients’ data share the same feature space. In this context, more similar label distributions imply more similar overall data distributions. These clients’ models can then serve as effective reference models for performing LiRA against the target client (Shokri et al., 2017; Carlini et al., 2022). While we notice that when target samples appear in other clients’ private datasets, those clients’ models cannot serve as reference models. However, this limitation is minimized in real-world FD deployments where clients typically have mostly disjoint training datasets to benefit from collaborative learning.